# Natural Language Processing for the Legal Domain: A Survey of Tasks, Datasets, Models, and Challenges

### v2

@@ 3.1.2. Domains with Shared Characteristics
Other domains exhibit features similar to those of legal texts, such as specialised vocabularies, large-scale corpora, and cross-referencing. In the medical domain, texts—including clinical notes, patient records, and research articles—rely on domain-specific terminology, diagnostic labels, and biological taxonomies, and often involve case-based analyses akin to legal reasoning. Similarly, software documentation and source code contain specialised programming terms and algorithmic descriptions, and are characterised by large repositories with numerous cross-references to libraries and functions.
@@ 3.2. Legal NLP
3.2.1. Introduction to legal NLP
Legal NLP involves the application of NLP techniques to legal texts. This field is crucial as it can help automate and enhance the analysis of complex legal documents, compliance, and decision-making processes (Huang et al., 2024). The foundation of NLP is text, and the legal domain primarily consists of textual data (Almuslim and Inkpen, 2022), including statutes, case law, contracts, and regulations. Given the text-intensive nature of the legal field, NLP offers potential to change how legal professionals interact with and use this information. By leveraging advanced algorithms and Machine Learning (ML) models, legal NLP aims to make legal texts more accessible, interpretable, and actionable (Huang et al., 2024).
@@ 3.2.2. Basic foundations and concepts of NLP
The integration of NLP in the legal domain relies on foundational techniques that enable the processing and analysis of legal texts. These techniques form the basis for various applications, transforming unstructured legal documents into structured, actionable information. This section introduces key NLP methods, machine learning paradigms, and text retrieval technique.
@@
    • Tokenization: Tokenization is the process of breaking text into smaller units, typically words or sub-words, known as tokens. This is a fundamental step in NLP, allowing structured text analysis. In legal NLP, tokenization can facilitate processing lengthy documents by segmenting them into manageable parts. Tokenization prepares the input text for numerical representation through word embeddings.
    • Word Embeddings:It represent words as continuous vectors in a high-dimensional space, capturing their semantic relationships. These embeddings enable models to encode word meanings and relationships, which are essential for tasks such as legal text similarity analysis and document classification. Word embeddings provide a numerical representation of tokens, which is essential for transformer-based models to process contextual relationships.
    • Transformers: The transformer architecture is a NN model designed to process text by first identifying connections between input words using self-attention. In this process, the model computes attention weights (which determine the importance of each word based on its context) for the entire input sequence. Only after establishing these connections does it generate output representations. Transformers operate on word embeddings and refine these embeddings by incorporating context from the entire sequence.
    • PLMs: PLMs, such as Bidirectional Encoder Representations from Transformers (BERT) (Devlin et al., 2019), are transformer-based models that have been trained on large-scale text corpora in a self-supervised manner. These models learn linguistic patterns and can be fine-tuned for specific legal NLP tasks. PLMs leverage transformer architectures and pre-trained embeddings to perform NLP tasks.
@@
    • Parameter-Efficient Fine-Tuning (PEFT): PEFT is a method for adapting PLMs that involves freezing the majority of the model’s parameters and only updating a small subset. This approach significantly reduces the computational resources and time required for fine-tuning, making it particularly effective in resource-limited scenarios, while still achieving competitive performance in tasks, such as text generation (Li et al., 2024a).
    These machine learning paradigms enhance the performance of core NLP methods, particularly PLMs, by tailoring models to specific tasks and optimising their training efficiency.
@@ 3.2.3. LLMs as an application of NLP in the legal sector
LLMs are a category of Deep Learning (DL)-based NLP models trained on extensive text corpora to process and generate human-like language. These models, typically based on Transformer architectures, learn linguistic patterns from diverse sources, enabling them to perform tasks, such as text summarisation, document analysis, and Question Answering (QA). LLMs, as part of NLP applications, have gained widespread attention, particularly with the release of ChatGPT in November 2022 (Ouyang et al., 2022). However, the legal sector has been exploring AI-driven solutions for much longer, applying NLP techniques to automate legal processes. NLP applications in the legal field include drafting client briefs, analysing large document sets, and enabling smaller firms to compete more effectively with larger ones (McGee, [n. d.]). NLP is very important in doing thorough checks when companies merge and greatly helps in legal education and learning in fast-changing fields (McGee, [n. d.]).
A notable demonstration of NLP’s potential in the legal sector occurred when GPT‑4 was reported to have passed the Uniform Bar Exam near the 90th percentile, underscoring the technology’s promise (Katz et al., 2024).However, subsequent analyses by Martínez (2024) suggest that its actual performance may be considerably lower, possibly around the 48th percentile overall and 15th percentile on essays. Meanwhile, further research reveals that although ChatGPT can achieve moderate success on certain legal classification tasks, smaller fine-tuned models still outperform it by about 30 percentage points (Chalkidis, 2023). Another study highlights a tendency for LLMs to hallucinate legal content at high rates—up to 58% in some cases—raising reliability concerns for complex tasks (Dahl et al., 2024). Despite these drawbacks, lawyers and law students remain conscious of the broader impact of such models: a recent LexisNexis survey (LexisNexis, [n. d.]) shows that about half of all lawyers believe LLMs will transform legal practice, with 92% anticipating some impact. Furthermore, 77% of respondents foresee efficiency gains for legal professionals, and 63% predict changes in how law is taught and studied.
@@ 3.2.4. Key publications and conferences in legal NLP
This section highlights the key journals, conferences, and workshops that serve as platforms for sharing advancements and insights in the intersection of NLP and the legal domain. These resources provide good opportunities for researchers to engage with cutting-edge research in legal NLP.
Several leading journals focus on the intersection of AI, NLP, and the legal domain. “Artificial Intelligence and Law”, published by Springer, is a leading journal that features research articles on legal reasoning, legal IR, and legal knowledge representation. A recent special issue, Applications and Evaluation of LLMs in the Legal Domain, examines the use of LLMs in legal tasks such as summarisation, judgement prediction, and contract drafting, while also addressing concerns, such as bias, misinformation, and regulatory compliance. Additionally, the “Journal of Law and Information Technology” focuses on the application of information technology in law, including research AI.
Conferences significantly advance research and promote collaboration in legal NLP. The International Conference on Artificial Intelligence and Law (ICAIL) is a biennial event that showcases advances in AI applications in the legal domain, including NLP and ML. The Conference on Legal Knowledge and Information Systems (JURIX) is an annual event that focuses on legal informatics and NLP technologies. In addition to these dedicated venues, prominent legal NLP research has also been published in broader AI, NLP, and IR conferences, including NeurIPS, ACL (and its associated workshops), NAACL, EMNLP, SIGIR, IJCAI, AAAI, and LREC.
In legal NLP, there are some good works which are presented in workshops. The workshop on Automated Semantic Analysis of Information in legal texts focuses on NLP and semantic analysis of legal texts. The international workshop on Juris-Informatics (JURISIN) brings together researchers from law, social science, and technology to discuss foundational and practical issues at the intersection of legal theory and informatics. The Natural Legal Language Processing (NLLP) workshop provides a platform for discussing NLP technologies tailored for legal texts and is often part of major NLP conferences. The EXplainable AI in Law (XAILA) workshop focuses on the explainability of AI systems in legal contexts, aiming to improve transparency and trust in AI applications. The Competition on Legal Information Extraction/Entailment (COLIEE) is an annual event that challenges participants to develop innovative solutions for legal information extraction and entailment tasks. Additionally, the Legal Track at the TREC, which ran from 2006 to 2011, focused on evaluating IR techniques for legal document review, providing a venue for researchers to test and compare methods in the context of e-discovery. Its contributions include benchmark datasets and evaluation metrics that were used to assess retrieval performance in legal text processing.
@@ 4. Methodology
This survey follows the Preferred Reporting Items for Systematic Reviews and Meta-Analyses (PRISMA) framework (Page et al., 2021). It ensures a transparent and comprehensive assessment of research to NLP tasks within the legal sector.
@@ 4.1. Search Strategy
We performed a systematic search across two academic databases to identify relevant studies, including: Google Scholar and IEEE Xplore. Then, search queries were crafted to capture studies that focused on the application of NLP to legal tasks. The search was defined by the following queries:
@@
Our search covered publications with the following date ranges for each specific NLP task: LQA from 2020-2024, LJP from 2017-2024, LTC from 2018-2023, LDS from 2016-2024, legal NER from 2010-2022, LAM from 2009-2024, legal corpora from 2021-2024, and legal LMs from 2020-2024. This ensured that we included recent advancements. Peer-reviewed journal articles and high-quality conference proceedings were prioritised, along with a secondary consideration of non-peer-reviewed sources, such as arXiv articles, where relevant.
@@ 4.2. Study Selection
A total of 154 studies were initially identified from the database search. To refine this list, we applied a manual reviewing and. This process involved:
@@
    Full-Text Review: Articles that passed the initial screening went through a more detailed full-text review. This was done to confirm their relevance, quality, and alignment with the inclusion criteria. During this phase, a careful study of the literature review sections of each included research and survey was conducted. This helped ensure that the studies not only contributed original findings but also reflected a comprehensive understanding of the existing research landscape in legal NLP.
@@
    Final Selection: From the original 154 studies, 133 were included in the final list, selected based on their direct relevance to key NLP tasks within the legal sector and their methodological quality, as well as their engagement with existing literature in the field.
@@ 4.3. Eligibility Criteria
    • The study must focus on at least one of the following NLP tasks: LQA, legal NER, LJP, LDS, LTC, LAM, legal LMs, or legal corpora.
@@
    • Studies focused exclusively on unrelated areas such as IR methods, pattern mining, information extraction, or similarity detection without a clear application to the specific legal NLP tasks mentioned.
@@
    • Editorials, opinion pieces, or other non-research articles.
@@ 5. Legal Question Answering
LQA involves responding to queries about law. This task is typically done by legal professionals skilled in the relevant legal field. This process requires a comprehensive review of existing laws, careful interpretation of statutes and regulations, and the application of legal principles and past cases to the relevant facts. LQA aims to provide advice on legal matters. It helps people and businesses in navigating the legal landscape and addressing legal challenges.
5.1. Task
@@
Recently, DL has been leveraged in LQA to employ NN models that train on extensive datasets to identify complex patterns and relationships. These models evaluate the questions posed, identify relevant legal topics, and produce suitable answers based on the patterns they have learned.
@@
Modern ML approaches for LQA rely on NN to understand natural language. Popular architectures include Recurrent Neural Networks (RNN), Long Short-Term Memory (LSTM), and Convolutional Neural Networks (CNN), which can be fine-tuned for specific tasks such as QA. These models generate better responses, adapt to new patterns, understand context, and provide relevant answers. Transformer-based models, such as BERT and ChatGPT, have proven particularly effective in NLP tasks. These models use the transformer architecture and self-attention mechanisms to learn the context of the text and understand the meaning of words. This allows them to provide relevant answers by weighing the importance of different parts of the input. In the following, we will study the existing LQA works in legal domain.
@@
Huang et al. (2020) introduce the Artificial Intelligence Law Assistant, the first Chinese LQA system that integrates a legal Knowledge Graph (KG) to enhance query comprehension and answer ranking. The system collects a large-scale QA corpus from an online legal forum and constructs a legal KG with over 42,000 legal concepts. It employs a knowledge-enhanced interactive attention network using Bidirectional LSTM (Bi-LSTM) and co-attention mechanisms to enrich semantic representations of QA pairs with legal domain knowledge. Additionally, it provides visual explanations for selected answers, offering users a clear understanding of the QA process.
An example of an answer retrieval system specific to Private International Law is proposed by Sovrano et al. (2020). This system integrates Term Frequency Inverse Document Frequency (TF-IDF) with deep LMs to retrieve relevant answers from an automatically generated KG of contextualised grammatical sub-trees. The KG aligns with a legal ontology based on Ontology Design Patterns, such as agent, role, event, temporal parameter, and action, to capture the relationships within and between provisions.
@@
Khazaeli et al. (2021) develop an IR-based QA system tailored to the legal domain, combining sparse vector search (BM25) and dense vector techniques (semantic embeddings) as input to a BERT-based answer re-ranking system. The system utilises Legal GloVe and Legal Siamese BERT embeddings to enhance retrieval performance. An answer finder component computes the probability of a passage answering the question using a BERT sequence binary classifier fine-tuned on question-answer pairs, improving the model’s ability to discriminate good answers.
@@
Li et al. (2024b) introduce a retrieving-then-answering framework featuring the Graph-Based Evidence Retrieval and Aggregation Network (GESAN) to enhance QA on the Judicial Examination of Chinese Question Answering (JEC-QA) dataset (Zhong et al., 2020c). The framework leverages relevant legal knowledge by predicting question topics and retrieving legal paragraphs using BM25. GESAN aggregates the evidence and processes it along with the question and options to make accurate predictions, demonstrating improved reasoning abilities in LQA.
Askari et al. (2022) propose a method for generating query-dependent textual profiles for lawyers to improve legal expert finding on QA platforms. Using data from the Avvo [^2] https://www.avvo.com QA forum, they focus on aspects such as sentiment, comments, and recency to create profiles. These profiles are fine-tuned using BERT, and the final aggregated score is calculated using a linear combination of profile-trained BERT. It improves retrieval performance in the legal expert finding task.
Zhang et al. (2023c) propose a generation-based method for LQA, modelling it as a generation task to produce new, relevant answers tailored to each question. The system incorporates laws as external knowledge into the answer generation process, using a retriever to fetch applicable law articles and a generator to create answers using this knowledge. Both components are integrated into a single T5 (Raffel et al., 2020) model using MTL. It can improve the model’s understanding and generation abilities while ensuring answers are accurate and informative.
Louis et al. (2024) present an end-to-end methodology for generating long-form answers to statutory law questions using a “retrieve-then-read” pipeline. The approach involves a retriever component that uses a bi-encoder model to fetch relevant legal provisions, followed by a generator that formulates comprehensive answers based on these provisions. The generator, an autoregressive LLM based on the Transformer architecture, employs in-context learning and PEFT to generate detailed answers. The model’s interpretability is enhanced by an extractive rationale generation strategy, ensuring responses are accompanied by verifiable justifications.
@@
Sovrano et al. (2024) propose DiscoLQA, a discourse-based LQA system that focuses on important discourse elements, such as Elementary Discourse Units and Abstract Meaning Representations. This approach helps the answer retriever identify the most relevant parts of the discourse, enhancing retrieval accuracy. They introduce the Q4EU dataset, containing over 70 questions and 200 answers on six European norms, demonstrating improved performance in legal QA even without domain-specific training.

### v3

@@ 3.1.2. Domains with Shared Characteristics
Other domains exhibit features similar to those of legal texts, such as specialised vocabularies, large-scale corpora and cross-referencing. In the medical domain, texts – including clinical notes, patient records, and research articles – rely on domain-specific terminology, diagnostic labels, and biological taxonomies, often involving case-based analyses akin to legal reasoning. Similarly, software documentation and source code contain specialised programming terms and algorithmic descriptions and are characterised by large repositories with numerous cross-references to libraries and functions.
@@ 3.2. Legal NLP
The legal sector has been exploring AI-driven solutions since the late 20th century, applying NLP techniques to automate legal processes. NLP applications in the legal field include drafting client briefs and analysing large document sets, enabling smaller firms to compete more effectively with larger ones (McGee, [n. d.]). These applications are also very important for compliance and due-diligence checks – required when companies merge their business, for instance – and greatly supports legal education and learning in fast-changing fields (McGee, [n. d.]). Legal NLP can also enhance the analysis of complex legal documents, thereby assisting with complex decision-making processes (Huang et al., 2024).
The foundation of NLP is text and the legal domain primarily consists of textual data (Almuslim and Inkpen, 2022), including statutes, case law, contracts and regulations. Given the text-intensive nature of the legal field, NLP offers potential to change how legal professionals interact with and use this information. By leveraging advanced algorithms and Machine Learning (ML) models, legal NLP aims to make legal texts more accessible, interpretable and actionable (Huang et al., 2024).
@@ 3.2.1. Basic foundations and concepts of NLP
The integration of NLP in the legal domain relies on foundational techniques that enable the processing and analysis of legal texts. These techniques form the basis for various applications, transforming unstructured legal documents into structured, actionable information. This section introduces key NLP methods, ML paradigms and text retrieval technique.
@@
    • Tokenization: Tokenization is the process of breaking text into smaller units, typically words or subwords, known as tokens. It is a fundamental step in NLP, allowing structured or unstructured text to be analysed. In legal NLP, tokenization facilitates processing lengthy documents by segmenting them into manageable parts and prepares the input text for numerical representation through word embeddings.
    • Word Embeddings: Embeddings represent words as continuous (numeric) vectors in a high-dimensional space, designed to capture semantic relationships. These embeddings enable models to encode word meanings and relationships, which are essential for tasks such as legal text similarity analysis and document classification.
    • Transformers: The transformer architecture is a NN model designed to process text by first identifying connections between input words using self-attention. In this process, it computes attention weights (which determine the importance of each word based on its context) across the entire input sequence before generating output representations. Transformers operate on word embeddings and refine these embeddings by incorporating context from the entire sequence.
    • PLMs: PLMs, such as Bidirectional Encoder Representations from Transformers (BERT) (Devlin et al., 2019), are transformer-based models pre-trained on large text corpora using self-supervised objectives. These models learn linguistic patterns and can be fine-tuned for specific legal NLP tasks. PLMs leverage transformer architectures and pre-trained embeddings to perform NLP tasks.
@@
    • Parameter-Efficient Fine-Tuning (PEFT): PEFT is a method for adapting PLMs to new tasks that involves freezing the majority of the model’s parameters, relying on updating just a small subset “fine-tuned” to the downstream task. This approach significantly reduces the computational resources and time required for fine-tuning, making it particularly effective in resource-limited scenarios, while still achieving competitive performance in tasks, such as text generation (Li et al., 2024a).
    These ML paradigms enhance the performance of core NLP methods, particularly PLMs, by tailoring models to specific tasks and optimising their training efficiency.
@@ 3.2.2. LLMs as an application of NLP in the legal sector
LLMs are a category of Deep Learning (DL)-based NLP models trained on extensive text corpora to process and generate human-like language. These models, typically based on Transformer architectures, learn linguistic patterns from diverse sources, enabling them to perform tasks, such as text summarisation, document analysis and Question Answering (QA). LLMs have gained widespread attention in NLP applications, particularly following the release of ChatGPT in November 2022 (Ouyang et al., 2022).
A notable demonstration of NLP’s potential in the legal sector occurred when GPT‑4 was reported to have passed the Uniform Bar Exam near the 90th percentile, underscoring the technology’s potential (Katz et al., 2024). However, subsequent analyses by Mart´ınez (2024) suggest that its actual performance may be considerably lower, possibly around the 48th percentile overall and 15th percentile on essays. Similarly, further research reveals that although ChatGPT can achieve moderate success on certain legal classification tasks, smaller fine-tuned models still outperform it by about 30 percentage points (Chalkidis, 2023). Another study highlights a tendency for LLMs to hallucinate legal content at high rates—up to 58% in some cases, raising reliability concerns for complex tasks (Dahl et al., 2024). Despite these drawbacks, lawyers and law students remain conscious of the broader impact of such models: a recent LexisNexis survey (LexisNexis, [n. d.]) shows that about half of all lawyers believe LLMs will transform legal practice, with 92% anticipating at least some impact. Furthermore, 77% of respondents foresee efficiency gains for legal professionals and 63% predict changes in how law is taught and studied.
@@ 3.2.3. Key publications and conferences in legal NLP
This section highlights the key journals, conferences and workshops that serve as platforms for sharing advancements and insights at the intersection of NLP and the legal domain. These resources provide opportunities for researchers to engage with cutting-edge work in legal NLP.
Several leading journals focus on the intersection of AI, NLP and the legal domain. “Artificial Intelligence and Law,” published by Springer, is a leading journal that features research articles on legal reasoning, legal IR and legal knowledge representation. A recent special issue, Applications and Evaluation of LLMs in the Legal Domain, examines the use of LLMs in legal tasks such as summarisation, judgement prediction and contract drafting, while also addressing concerns, such as bias, misinformation and regulatory compliance. Additionally, the “Journal of Law and Information Technology” focuses on the application of information technology in law, including research AI.
Conferences significantly advance research and promote collaboration in legal NLP. The International Conference on Artificial Intelligence and Law (ICAIL) is a biennial event showcasing advances in AI applications for the legal domain, including NLP and ML. The Conference on Legal Knowledge and Information Systems (JURIX) is an annual event that focuses on legal informatics and NLP technologies. In addition to these dedicated venues, prominent legal NLP research has also been published in broader AI, NLP and IR conferences, including NeurIPS, ACL (and its associated workshops), NAACL, EMNLP, SIGIR, IJCAI, AAAI and LREC.
In legal NLP, workshops also attract strong research contributions. The workshop on Automated Semantic Analysis of Information in Legal Texts focuses on NLP and semantic analysis of legal documents. The International workshop on Juris-Informatics (JURISIN) brings together researchers from law, social science and technology to discuss foundational and practical issues at the intersection of legal theory and informatics. The Natural Legal Language Processing (NLLP) workshop provides a platform for discussing NLP technologies tailored for legal texts and is often part of major NLP conferences. The EXplainable AI in Law (XAILA) workshop focuses on the explainability of AI systems in legal contexts, aiming to improve transparency and trust in AI applications. The Competition on Legal Information Extraction/Entailment (COLIEE) is an annual event that challenges participants to develop innovative solutions for legal information extraction and entailment tasks. Additionally, the Legal Track at the Text Retrieval Conference (TREC), which ran from 2006 to 2011, focused on evaluating IR techniques for legal document review, providing a venue for researchers to test and compare methods in the context of e-discovery. Its contributions include benchmark datasets and evaluation metrics that were used to assess retrieval performance in legal text processing.
@@ 4. Methodology
This survey follows the Preferred Reporting Items for Systematic Reviews and Meta-Analyses (PRISMA) framework (Page et al., 2021). It ensures a transparent and comprehensive assessment of research on NLP tasks within the legal sector.
@@ 4.1. Search Strategy
We performed a systematic search across two academic databases to identify relevant studies, including: Google Scholar and IEEE Xplore. Then, search queries were crafted to capture studies that focused on the application of NLP to legal tasks. The search was defined by the following two queries:
@@
Our search covered publications within the following date ranges for each NLP task: LQA from 2020-2024, LJP from 2017-2024, LTC from 2018-2023, LDS from 2016-2024, legal NER from 2010-2022, and LAM from 2009-2024. Furthermore, we limited our search for legal corpora to 2021-2024 and legal LMs from 2020-2024. This approach ensured the inclusion of recent advancements. Peer-reviewed journal articles and high-quality conference proceedings were prioritised, with secondary consideration given to relevant non-peer-reviewed sources.
@@ 4.2. Study Selection
A total of 154 studies were initially identified from the database search. To refine this list, we applied manual review. This process involved:
@@
    Full-Text Review: Articles that passed the initial screening underwent a detailed full-text review to confirm their relevance, quality and alignment with the inclusion criteria. During this phase, we also examined the literature review sections of each paper to ensure that the studies not only contributed original findings but also demonstrated a comprehensive understanding of the existing legal NLP research landscape.
@@
    Final Selection: Of the original 154 studies, 131 met the inclusion criteria and were retained. These studies were selected for their direct relevance to key legal NLP tasks, methodological quality, and engagement with existing literature.
@@ 4.3. Eligibility Criteria
    • The study must focus on at least one of the target NLP tasks (LQA, legal NER, LJP, LDS, LTC, LAM), or it must focus on legal LMs or legal corpora.
@@
    • Studies focused exclusively on unrelated areas such as IR methods, pattern mining, information extraction or similarity detection without a clear application to the specific legal NLP tasks mentioned.
@@
    • Editorials, opinion pieces or other non-research articles.
@@ 5. Legal Question Answering
LQA involves responding to legal queries, a task typically performed by professionals with domain expertise. It requires a comprehensive review of relevant laws, careful interpretation of statutes and regulations, and the application of legal principles and precedent to the facts. LQA aims to provide legal advice, helping individuals and businesses navigate the complex legal landscape.
5.1. Datasets
LQA datasets are a specialised resource designed to facilitate research in the domain of legal NLP. They consist of a collection of legal questions and corresponding answers, drawn from various legal documents and case law. Most questions in the LQA datasets fall into two main categories: knowledge-driven questions (KD-questions) and case-analysis questions (CA-questions) (Zhong et al., 2020c). KD-questions are centred around the understanding of specific legal concepts, whereas CA-questions involve the analysis of actual legal cases. Both categories demand advanced reasoning skills and a deep comprehension of the text, making LQA a particularly challenging task in the field of NLP.
Zhong et al. (2020c) introduce JEC-QA, a dataset with 26,365 multiple-choice questions from the National Judicial Examination of China and related websites. Each question provides four possible answers and is annotated with the type of reasoning required, such as word matching, conceptual understanding, numerical analysis, multi-paragraph comprehension and multi-hop inference. This dataset poses challenges for QA models, highlighting the gap between machine performance and human expertise in legal reasoning.
Sovrano et al. (2021) present the Q4PIL dataset, designed to evaluate automated QA systems in the domain of Private International Law. It includes 17 carefully selected questions based on key EU regulations – Rome I, Rome II and Brussels I bis – with answers derived directly from these regulations. The questions are classified based on their specificity, allowing for nuanced analysis of context-dependency in legal reasoning. This dataset supports the assessment of QA systems intended for legal professionals navigating complex cross-border issues.
Büttner and Habernal (2024) introduce GerLayQA, a dataset supporting LQA for laypersons in Germany, focusing on the civil-law system. It contains 21,538 real-world questions posed by laypersons, paired with expert answers from lawyers grounded in specific paragraphs of German legal codes. The dataset was constructed through filtering and quality assurance to ensure accuracy and relevance, making it a valuable resource for developing LQA systems that interpret and apply German law to everyday legal inquiries.
5.2. Approaches
@@
Recently, DL has been applied in LQA through NN models trained on large datasets to identify complex patterns and relationships. These models analyse posed questions, recognise relevant legal topics, and generate appropriate answers based on learned patterns and memory.
@@
Modern ML approaches to LQA use NN architectures to process natural language. Popular architectures include Recurrent Neural Networks (RNN), Long Short-Term Memory (LSTM), and Convolutional Neural Networks (CNN), which can be fine-tuned for QA tasks. These models adapt to new patterns, capture contextual information and generate more accurate responses. Transformer-based models, such as BERT and, more recently, ChatGPT, have proven particularly effective in NLP tasks. These models use the transformer architecture and self-attention mechanisms to learn the context of text to model the patterns and word dependencies in text. This allows them to provide relevant answers by weighting the importance of different parts of the input based on its contextual importance. In the following paragraphs, we will study the existing LQA works in the legal domain.
@@
Huang et al. (2020) introduce the Artificial Intelligence Law Assistant, the first Chinese LQA system that integrates a legal Knowledge Graph (KG) to enhance query comprehension and answer ranking. The system collects a large-scale QA corpus from an online legal forum and constructs a legal KG with over 42,000 legal concepts. It employs a knowledge-enhanced interactive attention network using Bidirectional LSTM (Bi-LSTM) and co-attention mechanisms to enrich semantic representations of question–answer pairs with legal domain knowledge. Additionally, it provides visual explanations for selected answers, offering users a clear understanding of the QA process.
@@
Khazaeli et al. (2021) develop an IR-based QA system tailored to the legal domain, combining sparse term-based search (BM25) and dense vector techniques (semantic embeddings) as input to a BERT-based answer re-ranking module. The system utilises Legal GloVe and Legal Siamese BERT embeddings to enhance retrieval performance. An “answer finder” component computes the probability that a passage answers the question using a BERT sequence classifier fine-tuned on question–answer pairs, thereby enhancing the model’s ability to discriminate relevant answers.
@@
Li et al. (2024b) introduce a retrieve-then-answer framework featuring a Graph-Based Evidence Retrieval and Aggregation Network (GESAN) to enhance LQA on the JEC-QA dataset (Zhong et al., 2020c). The framework leverages legal knowledge by predicting question topics and retrieving relevant paragraphs using BM25. GESAN aggregates the evidence and processes it along with the question and options to generate accurate predictions, demonstrating improved reasoning capabilities in LQA.
Askari et al. (2022) tackle legal expert finding on QA platforms by building query-dependent textual profiles for lawyers. Using data from the Avvo forum, [^3] https://www.avvo.com they represent each lawyer with four facets – comment content, positive sentiment, negative sentiment, and answer recency – derived from past answers and comments. A separate BERT model is fine-tuned for each facet and the scores are linearly combined with a document-based BERT ranker to produce the final ranking.
[…]
Louis et al. (2024) propose an end-to-end “retrieve-then-read” system that generates long-form answers to statutory questions. A lightweight bi-encoder first retrieves the most relevant French legal provisions, after which an instruction-tuned LLM, adapted via in-context learning and PEFT, composes detailed answers that cite those statutes. To ensure transparency, the model also outputs extractive rationales, listing the exact paragraphs that justify each response.
@@
Sovrano et al. (2024) propose DiscoLQA, a discourse-based LQA system that focuses on important discourse elements, such as Elementary Discourse Units and Abstract Meaning Representations. This approach helps the answer retriever identify the most relevant parts of the discourse, enhancing retrieval accuracy. They introduce the Q4EU dataset, containing over 70 questions and 200 answers on six European norms, demonstrating improved performance in LQA even without domain-specific training.
@@
6. Legal Judgement Prediction
LJP is a key task in legal NLP, particularly in civil‐law jurisdictions where judgements rely on case facts and statutory provisions (Zhong et al., 2020b). It seeks to predict legal outcomes from case descriptions and the applicable legislation (Zhong et al., 2020b). The topic has attracted growing interest from AI researchers and legal professionals because of its potential to help judges, lawyers and scholars anticipate case results based on historical data (Chen et al., 2024a).
LJP is clearly a demanding and complex problem. Historical legal data can contain inherent biases that can create feedback loops and amplify discrimination if not managed carefully (Krasadakis et al., 2024). Therefore, ensuring the impartiality of predicted rulings is crucial (Krasadakis et al., 2024). At present, LJP is still performed chiefly by legal experts who require extensive specialised training to identify relevant statutes, define charge ranges, and set penalty terms (Cui et al., 2023).